OpenStudy (anonymous):

Find d/dxm(3x-5)/(x^2+7)

OpenStudy (anonymous):

what is the m doing there?

OpenStudy (anonymous):

Nothing it was a typo mistake

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

quotient rule for this one

OpenStudy (anonymous):

\[f(x) = \frac{ 3x-5 }{ x^2+7 }\]

OpenStudy (misty1212):

\[\left(\frac{f}{g}\right)'=\frac{gf'-fg'}{g^2}\]with \[f(x)=3x-5.f'(x)=3,g(x)=x^2+7,g,(x)=2x\] grind it till you find it